# Design Optimizer for Soft Growing Robot Manipulators in Three-Dimensional Environments

### v1

@@ 1 Introduction
Soft robotics is a relatively new field of intelligent mechanical systems that aims at overcoming some of the biggest limitations of their rigid counterparts [1]. Rigid robots (i) may struggle to interact with fragile objects or delicate environments, such as human tissues, glassware, or sensitive components in archaeological sites; (ii) have intricate designs and their materials often result in high costs; and (iii) due to their strength and stiffness, the safety of human operators can be a concern in a human-machine interaction scenario [2]. On the other hand, soft robots are inspired by living organisms, with designs exploiting the natural compliance of their materials to facilitate smooth and intuitive physical interactions with the environment [3, 4, 1, 5]. Moreover, their materials and actuation methods are typically more affordable, safer, and simpler to manufacture or assemble compared to conventional rigid robots [6]. Additionally, soft robots have the ability to extend or retract through a process known as eversion – i.e., like a linear actuator, but with the advantage of occupying no space when completely retracted because the soft material can be folded. This unique degree of freedom (DoF) is directly inspired by the growth patterns of plants, leading to these robots often being referred to as vine robots [7, 8, 9] – or more generically soft growing robots. Several studies have demonstrated their effectiveness in manipulation tasks that require payload-handling capabilities [10, 11, 12, 13]. A major innovation was the introduction of continuum links and discrete joints, which can be achieved using layer jamming to dynamically adjust the robot’s stiffness at the joint locations to independently steer different links in different directions [14]. As illustrated in Fig. 1(a), this allows soft growing robots to be modeled with the classic kinematic chain used for their rigid counterparts, with the substantial difference that each link also grows and retracts.
@@
Figure 1: (a) A soft growing manipulator with continuum links, variable stiffness, and variable discrete joints [14]. (b) Sketch of a three-dimensional soft growing robot reaching two targets in a maze (side view). The robot progressively grows, retracts, and steers to reach the next target. (c) A real prototype of a soft growing manipulator.
@@
Soft robots are considerably less costly than their rigid counterparts, and this makes them accessible to the general public, including non-professional robot designers. Virtually anyone can create their own soft growing manipulator for their specific tasks [15, 16, 17]. However, specifying and integrating the components needed for a task-specific robot remains a complex challenge for human designers. With the capability of growing and retracting, the robot’s tip (namely the end effector or where the gripper/manipulative device is placed) can move anywhere along its kinematic chain. Consequently, their kinematic modeling and actuation pose significant difficulties. Ideally, a soft-robot designer would only need to define the functional goals and constraints of a given task, and an optimal design would be automatically generated [18].
@@
Determining the design of a soft growing robot is a complex engineering problem that can be addressed using mathematical optimization [19]. A design defines the parameters and attributes describing a soft growing robot, ultimately guiding its manufacturing process. This includes aspects such as link lengths, joint positions and their range of motion (if applicable), materials used, and the overall shape or type of the robot. This problem has been addressed with several optimization methods: the interior point algorithm [20, 21, 22, 23, 24, 25], the Levenberg-Marquardt algorithm [20, 26], simplex methods [27, 28, 29, 30, 31, 32], the steepest descent method [33, 34], linear approximations [35], cyclic coordinate descent [36], Bayesian optimization [25], the method of moving asymptotes [37, 38], the Newton-Raphson method [39], greedy algorithms [40], hill climbing [41], simulated annealing [25], pattern search [42, 43], specific heuristics [44, 45], the parametric quadratic programming algorithm [46, 47], and the Broyden-Fletcher-Goldfarb-Shanno algorithm [48]. However, these traditional optimization methods often struggle with navigating highly complex, nonlinear, and multimodal design spaces, making it difficult to find global optima in engineering design problems. That is why engineering problems often rely on artificial intelligence, specifically the subfield of Evolutionary Computation (EC), a sub-field of Artificial Intelligence [49]. EC algorithms, which are natural-inspired metaheuristic optimizers, are widely used in mechanical and robotic design [50, 51, 52, 53, 54, 55]. Specifically for soft robotics, many studies have used EC algorithms to solve their design optimization problems: genetic algorithms for single [56, 18, 57, 58, 59, 60, 27, 61, 41] and multi-objective optimization [62, 63, 64, 65, 66, 67], particle swarm optimization [27, 68, 69, 70, 71], differential evolution [20], age-fitness-pareto optimization [72], the estimation of distribution algorithm [73], covariance matrix adaptation evolution strategy [61, 74, 75], and neuroevolution [65, 76]. These studies aim to optimize a soft robot’s design by finding the simplest possible configuration (i.e., the minimum number of sections) capable of performing a specific task or set of tasks [43]. The design optimization typically requires solving the robot’s kinematics, which, in the case of a soft manipulator, also involves optimizing its joint bending angles or curvature at each point in time [70].
@@ 2 Soft Growing Manipulator for the Proposed Scenario
The soft growing manipulator examined in this study is based on the design introduced by Do et al. [14], which combines characteristics of both rigid and soft robots – Fig. 2 depicts a portion of the manipulator’s model to show its DoFs. The manipulator consists of links that can be individually adjusted to stiffen or stay flexible. When a link is not stiffened, it can be deliberately steered and maneuvered using cables to function as a revolute joint. When a link is stiffened, pulling its respective cable will not change its orientation; however, the following link (if not stiffeners) will be steered in the desired direction. Thanks to this technology, each link can steer independently by using only two cables for each direction (i.e., clockwise and anti-clockwise rotation around the x-axis, clockwise and anti-clockwise rotation around the y-axis). Lastly, growth and retraction (eversion) are made possible by the structure of the robot’s body, which is an inflatable beam composed of a thin-walled, compliant material that everts pneumatically. Starting in a fully folded state, pressurizing the interior with air causes the robot to unfold at its tip, pulling new material through its body [8, 79, 80].
@@
3 Optimization Problem and Proposed Solution
The problem requires us to find the best design for a soft growing robot manipulator to solve a specific task – i.e., the set of link lengths composing the body of the robot such that the robot can easily reach every target in the environment. Our optimizer will retrieve the design automatically. This problem can be defined in both 2D and 3D based on the effect of gravity on the system: a robot exploring a tunnel or a pipe system can be thought of as a 2D robot, but when the environment demands that the robot explore different levels (e.g., a multi-story building), it must also navigate in the third dimension. The 2D scenario was tackled in our previous work [67], and the current work extends it to the third dimension. Fig. 4 shows a 3D representation of the problem, where the robot is mounted on the environment’s ceiling and grows in favor of gravity [^1] In this work, gravity is oriented in the positive direction of the z-axis.. To simplify the problem, the model of our robot does not have gravity constraints on its orientation and can steer in any direction. Fig. 4(a) shows an example of a 3D task, in which the manipulator needs to reach three targets with a specified reaching orientation: this requires the robot to align its last link(s) to a specific straight line. Fig. 4(b) shows a possible solution to the problem, with three configurations (one for each target) sharing the same link lengths except for the last one.
@@ 3 Optimization Problem and Proposed Solution
Figure 4: (a) Example of a task to be performed by the manipulator: three targets placed on different pillars with a given orientation. Starting from the home base $h$, the manipulator will grow such that its final link is aligned with each target. This task has three targets ($|t|=3$) and three obstacles ($|o|=3$) – i.e., the pillars where the targets are placed. (b) Example of solution $\psi$ provided by the optimizer. The solution $\psi$ comprises three configurations $\zeta_{1\rightarrow 3}$ sharing the same design $\delta$ (i.e., the lengths of links are shared among the three configurations). Each configuration can have a different number of joints, as the robot will employ only a sufficient number of links to reach a certain target. In this solution, the links whose endpoint corresponds to the tip of the robot are $\bar{n}_{1}=4$, $\bar{n}_{2}=5$, $\bar{n}_{3}=4$.
The robot begins its extension from its container, known as the home base. Based on the home base’s pose (position and orientation) and the set of targets to be reached, the optimizer determines an optimal robot design that allows it to reach all targets while satisfying hardware constraints (see Section 3.5). The design parameters include the number of joints and their placement along the robot: these parameters must be determined prior to the robot’s manufacturing and assembly – specifically, the lengths of the robot’s links. However, the number of joints or links that a soft growing robot can utilize during a task may change depending on the target it needs to reach. Additionally, the length of the final link may vary, as it depends on how much the robot has extended to reach a single target. Unlike rigid robots, where the design space is simply defined by the length of each link, the design space for a soft growing robot has unique characteristics due to its ability to grow and retract: (i) different targets may require varying numbers of links to be reached; (ii) interior links have a fixed length across different targets, while the length of the final link can change for each target but cannot exceed its maximum length (which equals to the length it would have if it were an interior link); (iii) a link can serve as either an interior or a final link for different targets; and (iv) the number of links required for a given set of targets is not predetermined and changes based on the task. Fig. 4(b) shows a single robot with fixed link lengths designed to reach three targets with a given pose: each robot configuration $\zeta_{i}$ (one for each target) has the same link lengths, except for the final one, as the specific target $t_{i}$ might not require additional extension, leaving the remainder of the robot folded within its body.
@@
We model this scenario as an optimization problem: beginning with a random solution (random link lengths and joint angles for each configuration), we aim to minimize the distance between the robot’s tip and the targets’ poses using forward kinematics. Hence, the optimizer works as an inverse kinematics solver, ultimately determining the set of angles and link lengths (design) needed to reach each target.
3.1 Problem Formulation and Representation of the Decision Variables
The mathematical formulation of the problem defines the following terms:
@@ 3.1 Problem Formulation and Representation of the Decision Variables
    $n$, the maximum number of links;
@@
    $\theta^{(x)}$, joint rotation around x-axis;
@@
    $\theta^{(y)}$, joint rotation around y-axis; and
  • 4.
    $l$, link length.
Any start point of a link is defined as a node, which also corresponds to a set of two rotational joints (around x and y) and a linear joint (for growth and retraction). Only the last endpoint (i.e., the end effector or the node at the end of the last link) is not considered a joint set but simply the tip of the robot – i.e., not a node. Therefore, a robot will have at most $n$ nodes plus the end effector.
The user will set the bounds on these parameters based on the resources available for manufacturing – which might depend on budget, material availability, time, or the size of the machinery used for production. The number of nodes $n$ is defined in $\mathbb{N}^{+}$ and depends on how many links the manufacturer is prepared to produce. Angles on x and y are constrained within a range $\Delta_{\theta}\in\mathbb{R}$, which is determined by hardware limitations (e.g., each rotational joint cannot rotate more than $\pi/4$); this excludes the first joint at the base, where the robot extends orthogonally (i.e., either fixed at zero or with an unrestricted range if the home base is manually oriented for each target). Link lengths are constrained within a range $\Delta_{l}\in\mathbb{R}$ and are primarily influenced by manufacturing limitations (e.g., if the links are made from polyethylene plastic using a laser cutter, the maximum length is determined by the cutter’s size). Additionally, they must accommodate the gripping mechanism, meaning the link length must be at least as long as the gripper.
A specific task is defined with the following:
@@
    $t$, a set of targets, each represented by their pose in the space (three coordinates for position and two angles for desired reaching orientation);
@@
    $o$, a set of obstacles, each represented by their position in the space and other features – defined based on their geometry [^2] As described in Section 3.5, we represent all obstacles as cylinders oriented orthogonally with the floor, therefore the features of an obstacle include their radius, their height, and the coordinates of its origin as the center of the base.; and
@@
    $h$, the pose of the home base (three coordinates for position and two angles for orientation) – although by default, we always place the home base to the origin of the reference frame, oriented toward the positive direction of the z-axis.
The result of the optimization problem consists of a design $\delta$, a solution $\psi$, and a set of $t$ configurations $\zeta_{i}$ (one for each target $i=1...|t|$):
@@
    a single configuration $\zeta$ is defined as in Eqn. (1), and it is composed by the steering angles over x and y-axes at each joint and the length of each link;
@@
    multiple configurations define a solution [^3] A solution to an optimization problem is a well-defined set of values for the decision variables; this is also known as an individual in Evolutionary Computation. of the problem $\psi$, defined as in Eqn. (2) – note that each pair of rows represent steering angles on x and y, respectively, with the last row being reserved for shared link lengths (resulting in a $(2\cdot|t|+1)\times n$ matrix); and
@@
    the desired design for the specific task is defined by the vector $\delta$ as in Eqn. (3), which contains the values of the link lengths of the robot (from the first to the last) – this is the last row of the matrix $\psi$ since all configurations share the same design.
@@
3.2 Search Space
A higher number of links enhances the robot’s dexterity and expands its workspace. Consequently, the optimization problem must handle a large search space $\omega$, which grows linearly with the number of targets $|t|$, as it needs to optimize a single set of links for all targets. The search space $\omega$ is defined as in Eqn. (4) when the order of targets is predetermined, or if the robot returns to its base after reaching each target (i.e., the order of targets is irrelevant).
$\omega=n^{\Delta_{l}}+2\cdot|t|\cdot n^{\Delta_{\theta}}$ (4)
When the order of targets is unknown, and the solution must also minimize the cost of transitioning between configurations, then the search space grows factorially, as every possible permutation of target sequences must be considered. In this work, we do not address the target order and focus solely on solving the basic optimization problem. The order can be considered after optimization with a proper motion planner like the one proposed by Altagiuri et al. [83] to determine the most efficient target-reaching sequence.
3.3 Fitness Evaluation from the Objective Functions
The optimization problem is composed of the following objectives:
  1. 1.
    $f_{1}(\psi,t,h,o)$: minimize the distance between the end effector (which can be located anywhere along the body due to eversion) and each target in the environment – solve the inverse kinematics;
  2. 2.

### v2

@@ 1 Introduction
Soft robotics is a relatively new field of intelligent mechanical systems that aims at overcoming some of the biggest limitations of their rigid counterparts [1]. Rigid robots (i) may struggle to interact with delicate objects or environments, such as human tissues, glass objects, or sensitive components in archaeological sites; (ii) have intricate designs and their materials often result in high costs; and (iii) due to their strength and stiffness, the safety of human operators can be a concern in a human-machine interaction scenario [2]. On the other hand, soft robots are inspired by living organisms, with designs exploiting the natural compliance of their materials to facilitate smooth and intuitive physical interactions with the environment [3, 4, 1, 5]. Moreover, their materials and actuation methods are typically safer, cheaper, and simpler to manufacture or assemble compared to conventional rigid robots [6]. Additionally, soft robots have the ability to extend or retract through a process known as eversion – i.e., like a linear actuator, but with the advantage of occupying no space when completely retracted because the soft material can be folded. This unique degree of freedom (DoF) is inspired by the growth patterns of plants, leading to these robots often being referred to as vine robots [7, 8, 9] – or more generically soft growing robots. Several studies have demonstrated their effectiveness in manipulation tasks that require handling of payloads [10, 11, 12, 13]. A major innovation was introduced by Do et al. [14] with continuum links and discrete joints: thanks to dynamic layer jamming, the robot can independently stiffen the different parts of its body, resulting in a joint. As illustrated in Fig. 1(a), this allows soft growing robots to be modeled with the classic kinematic chain used for their rigid counterparts, with the substantial difference that each link also grows and retracts.
@@
Figure 1: (a) A soft growing manipulator with continuum links and variable discrete joints [14]. (b) Sketch of a three-dimensional soft growing robot reaching two targets in an environment with obstacles (side view). (c) A picture of a physical manipulator.
@@
Soft robots are considerably less costly than their rigid counterparts, and this makes them accessible to everyone, including amateur roboticists. Every robotic enthusiast could create their soft growing manipulator for one or more well-defined specific tasks [15, 16, 17]. However, specifying and integrating the components needed for a task-specific robot remains a complex challenge for human designers. With the capability of growing and retracting, the robot’s tip (namely the end effector or where the gripper/manipulative device is placed) can move anywhere along its kinematic chain. Consequently, their kinematic modeling and actuation pose significant difficulties. Preferably, a designer only needs to define the task itself, including any hardware or environmental constraint, and use an optimizer to retrieve optimal designs [18].
@@
Determining a soft growing robot’s design is a complex engineering problem that mathematical optimization can easily address [19]. A design defines the parameters and attributes describing a soft growing robot, ultimately guiding its manufacturing process. This includes aspects such as link lengths, joint positions and their range of motion (if applicable), materials used, and the overall shape or type of the robot. This problem has been addressed with several optimization methods: the interior point algorithm [20, 21, 22, 23, 24, 25], the Levenberg-Marquardt algorithm [20, 26], simplex methods [27, 28, 29, 30, 31, 32], the steepest descent method [33, 34], linear approximations [35], cyclic coordinate descent [36], Bayesian optimization [25], the method of moving asymptotes [37, 38], the Newton-Raphson method [39], greedy algorithms [40], hill climbing [41], simulated annealing [25], pattern search [42, 43], specific heuristics [44, 45], the parametric quadratic programming algorithm [46, 47], and the Broyden-Fletcher-Goldfarb-Shanno algorithm [48]. However, these traditional optimization methods often struggle with navigating highly complex, nonlinear, and multimodal design spaces, making it difficult to find global optima in engineering design problems. That is why engineering problems often rely on artificial intelligence, specifically the subfield of Evolutionary Computation (EC), a sub-field of Artificial Intelligence [49]. EC algorithms, which are natural-inspired metaheuristic optimizers, are widely used in mechanical and robotic design [50, 51, 52, 53, 54, 55]. Specifically for soft robotics, many studies have used EC algorithms to solve their design optimization problems: genetic algorithms for single [56, 18, 57, 58, 59, 60, 27, 61, 41] and multi-objective optimization [62, 63, 64, 65, 66, 67], particle swarm optimization [27, 68, 69, 70, 71], differential evolution [20], age-fitness-pareto optimization [72], the estimation of distribution algorithm [73], covariance matrix adaptation evolution strategy [61, 74, 75], and neuroevolution [65, 76]. These studies aim to optimize a soft robot’s design by finding the simplest possible configuration (i.e., the minimum number of sections) capable of performing a specific task or set of tasks [43]. The design optimization typically requires solving the robot’s kinematics, which, in the case of a soft manipulator, also involves finding the correct curvature along the body [70].
@@ 2 Soft Growing Manipulator for the Proposed Scenario
The soft growing manipulator examined in this study is based on the design proposed by Do et al. [14], which combines characteristics of both rigid and soft robots – Fig. 2 depicts a portion of the manipulator’s model to show its DoFs. The manipulator consists of links that can be individually adjusted to stiffen or stay flexible. When a link is not stiffened, it can be deliberately steered and maneuvered using cables to function as a revolute joint. When a link is stiffened, pulling its respective cable will not change its orientation; however, the following link (if not stiffeners) will be steered in the desired direction. Thanks to this technology, each link can steer independently by using only two cables for each direction (i.e., clockwise and anti-clockwise rotation around the x-axis, clockwise and anti-clockwise rotation around the y-axis). Lastly, growth and retraction (eversion) are made possible by the structure of the robot’s body, which is an inflatable beam composed of a thin-walled, compliant material that everts pneumatically. Starting in a fully folded state, pressurizing the interior with air unfolding the tip, pulling new material through its body [8, 79, 80].
@@
Figure 4: (a) Task example: three targets placed on different pillars with a desired orientation. The robot grows from its base $h$ such that its final link is aligned with each target. This task has three targets ($|t|=3$) and three obstacles ($|o|=3$) – i.e., the pillars where the targets are placed. (b) Optimizer’s output example: the solution $\psi$ comprises three configurations $\zeta_{1\rightarrow 3}$ that share a unique design $\delta$ (i.e., the three configurations $\zeta$ share the same link lengths). Because the robot may use only a limited amount of links to reach a specific target, depending on its distance to it, each configuration can have a different number of employed links. In this example, the number of links whose endpoint corresponds to the robot’s tip are $\bar{n}_{1}=4$, $\bar{n}_{2}=5$, $\bar{n}_{3}=4$.
3 Optimization Problem
This section describes the optimization problem, its formalization, and how the objectives were engineered to formulate a proper fitness evaluation. Since the details are similar to the 2D case, with the exception of adding a further dimension for the rotational joints, we briefly summarize this description and invite readers to find more details in our previous work [67]. However, unlike our previous work, more information has been added regarding the specifics of the procedures, which can be found in Section 1.1 of the Supplementary Material.
The problem requires us to find the best design for a soft growing robot manipulator to solve a specific task – i.e., the set of link lengths composing the body of the robot such that the robot can easily reach every target in the environment. Our optimizer will retrieve the design automatically. We defined this problem in both 2D and 3D based on the effect of gravity on the system: a robot navigating a tunnel or a pipe system can be thought of as 2D; but when the environment demands navigating different levels or layers, it requires 3D exploration. The 2D scenario was tackled in our previous work [67], and the current work extends it to the third dimension. Fig. 4 shows a 3D representation of the problem, where the robot is mounted on the environment’s ceiling and grows in favor of gravity – oriented in the positive direction of the z-axis. To simplify the problem, we did not consider gravity in the model. Fig. 4(a) shows an example of a 3D task in which the manipulator needs to reach three targets according to a desired reaching orientation. The robot should then align its last link – or more than one – so that it lies on the requested straight line. One example of design is shown in Fig. 4(b), with three configurations (one for each target) sharing the same link lengths except for the last one.
The robot begins its extension from its container, known as the home base $h$. Based on the home base’s pose (position and orientation) and the environment, the optimizer determines the optimal robot design so that the manipulator can reach every target without violating any constraint (see Section 3.3). The joint number and placement define the parameters of the design, which must be known prior to the robot’s manufacturing and assembly – specifically, the design $\delta$ is a set of link lengths defining a single manipulator.
This results in a single optimal solution $\psi$ having multiple configurations $\zeta$ for each target (i.e., a different manipulator’s orientation reaching each target) sharing a unique link-length design $\delta$. Configurations can differ from each other due to eversion: they may use a different link number because not all the links may actually be employed – i.e., they are folded within the robot’s body. Additionally, it is also not guaranteed that the last link will be completely everted, depending on how far the last employed joint is from the specific target.
@@ 3 Optimization Problem
We model this scenario as an optimization problem: beginning with a random solution (random link lengths and joint angles for each target/configuration) generated through forward kinematics, it minimizes the distance from the end effector and each of the targets’ poses. The design (link lengths and, consequently, link angles) are retrieved as an inverse kinematic solver.
Table 1: Nomenclature of the terms defined in the Optimization Problem
Symbol Description
$\delta$ design
$\psi$ solution
$\zeta$ configuration
$n$ maximum number of links/nodes
$\theta^{(x)}$ joint rotation around x-axis
$\theta^{(y)}$ joint rotation around y-axis
$l$ length of link
$t$ set of targets
$o$ set of obstacles
$h$ home base pose
$\epsilon$ index of the closest node to target’s orientation segment
$\theta^{(x)}_{\epsilon}$ joint rotation around x-axis at the $\epsilon$-link toward target
$\theta^{(y)}_{\epsilon}$ joint rotation around y-axis at the $\epsilon$-link toward target
$\bar{n}$ number of active/folded links
$l_{\bar{n}}$ length of the last active/folded link
$f_{1-2}$ objective function: inverse kinematics
$f_{3.1a}$ objective function: number of links to segment
$f_{3.1b}$ objective function: number of links on segment
$f_{3.2}$ objective function: overall robot length
$f_{3.3}$ objective function: undulation
The optimization problem is defined with the terminology summarized in Table 1.
3.1 Problem Formulation
We begin the formulation with the following:
@@ 3.1 Problem Formulation
    $n\in\mathbb{N}^{+}$, maximum link number, corresponding to the number of links the user is prepared to produce;
@@
    $\theta^{(x)}$ and $\theta^{(y)}\in\mathbb{R}$, rotation of joint around x and y-axes, respectively, and constrained within a domain $\Delta_{\theta}$, which is determined by how much the robot can bend without breaking or buckling – note that we decided to exclude the joint corresponding to the home base (i.e., the first) due to hardware observations so that the robot will grow from the base in the direction of its orientation and steer only from the second joint on; and
@@
    $l\in\mathbb{R}$, length of a link, constrained within a domain $\Delta_{l}$, and similar to the maximum link number $n$, it is influenced by production limitations (i.e., the overall size of the material sheet used for the body, the size of the cutting tool, etc.). Additionally, they have to be longer than the gripping mechanism to ensure stability on the end effector – i.e., imagine the gripper that moves along the body as the robot everts, pushed by the end effector [81]; this mechanism has a specific length, which should be smaller than any link length because if a joint happens to be inside the gripper, that joint cannot bend.
Any start point of a link is defined as a node, which also corresponds to a set of two rotational joints (around x and y) and a linear joint (for growth and retraction). Only the last endpoint (i.e., the end effector) is not considered a joint set but simply the robot’s tip – i.e., not a node. A design will have, at most, $n$ nodes plus the end effector.
The problem formulation also includes the task, which is defined as:
@@
    $t$, target set, defined as an array of poses (i.e., for each target, three coordinates for position and two angles for desired reaching orientation);
@@
    $o$, obstacle set, defined as an array of poses and geometrical properties [^1] In this work, obstacles are modeled as cylinders oriented orthogonally with the floor, described by a radius, height, and the coordinates of its origin as the center of the base – see in Section 3.3.; and
@@
    $h$, home base pose (three coordinates for position and two angles for orientation) – although by default, we always place the home base to the origin of the reference frame, oriented toward the positive direction of the z-axis.
The output of the optimization consists of the following terms:
@@
    configurations $\zeta_{i}$ (one for each target $i=1...|t|$), containing a column vector for the steering angles over the x-axis, y-axis, and the length of each link as in Eqn. (1);
@@
    a solution $\psi$ contains $|t|$ configurations, which provides the values of the decision variables and a description of the individuals’ genotype, defined as in Eqn. (2) – note that each pair of rows represent steering angles on x and y, respectively, with the last row being reserved for shared link lengths (resulting in a $(2\cdot|t|+1)\times n$ matrix); and
@@
    the final design $\delta$ defined as in Eqn. (3), containing link lengths as they appear in $\psi$ – i.e., one design for multiple configurations in the solution.
@@
3.2 Fitness Evaluation and Dynamic Genotype Change
We included the following objective functions:
  • 1.
    $f_{1}$: For each configuration, minimize the distance between the end effector and the target – inverse kinematic solver.
  • 2.